The total number of computers for Italy and Russia is 56,280 million. If Italy has 2340 million more computers than Russia, find the number of computers for each country.
This one really stumped me.
Thank you!!

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39627)

I = how many in Italy
R = how many in Russia

I+R=56280.
I=2340+R.
Two equations, two unknowns. Simple.

Answer by stanbon(75887)

The total number of computers for Italy and Russia is 56,280 million. If Italy has 2340 million more computers than Russia, find the number of computers for each country.
----
Equations:
I + R = 56,280 mil
I - R = 2340 mil
----
Add and solve for "I":
2*I = 58,620
---
I = 29310 million (# Italy has)
---
R = 56280-29310 = 26970 million (# Russia has)
=====================
Cheers,
Stan H.
================